# mddefDefinition

### v2

@@ Abstract
Machine learning force fields (MLFFs) are an attractive alternative to ab-initio methods for molecular dynamics (MD) simulations. However, they can produce unstable simulations, limiting their ability to model phenomena occurring over longer timescales and compromising the quality of estimated observables. To address these challenges, we present Stability-Aware Boltzmann Estimator (StABlE) Training, a multi-modal training procedure which leverages joint supervision from reference quantum-mechanical calculations and system observables. StABlE Training iteratively runs many MD simulations in parallel to seek out unstable regions, and corrects the instabilities via supervision with a reference observable. We achieve efficient end-to-end automatic differentiation through MD simulations using our Boltzmann Estimator, a generalization of implicit differentiation techniques to a broader class of stochastic algorithms. Unlike existing techniques based on active learning, our approach requires no additional ab-initio energy and forces calculations to correct instabilities. We demonstrate our methodology across organic molecules, tetrapeptides, and condensed phase systems, using three modern MLFF architectures. StABlE-trained models achieve significant improvements in simulation stability, data efficiency, and agreement with reference observables. By incorporating observables into the training process alongside first-principles calculations, StABlE Training can be viewed as a general semi-empirical framework applicable across MLFF architectures and systems. This makes it a powerful tool for training stable and accurate MLFFs, particularly in the absence of large reference datasets.
@@ 1 Introduction
Molecular dynamics (MD) simulation is a staple method of computational science, enabling high-resolution spatiotemporal modeling of atomistic systems throughout biology, chemistry, and materials science Frenkel & Smit (2001).While the atomic forces needed for MD simulation can be obtained on-the-fly via ab-initio quantum-mechanical (QM) calculations Car & Parrinello (1985), this is prohibitively expensive for realistic length and time scales Friesner (2005). Machine learning force fields (MLFFs) have recently emerged as a promising option to serve as surrogate models for QM calculations, demonstrating the ability to capture complex many-body interactions, and in some cases transfer flexibly across chemical space Schütt et al. (2018); Hu et al. (2021); Liu et al. (2022); Gasteiger et al. (2021; 2020); Batzner et al. (2022); Musaelian et al. (2022); Batatia et al. (2022); Schütt et al. (2021). Graph neural network (GNN)-based MLFFs trained on large ab-initio datasets are increasingly being used to model challenging and important chemical systems with favorable results Merchant et al. (2023); Kozinsky et al. (2023); Chen & Ong (2022); Schaarschmidt et al. (2022); Takamoto et al. (2022); Majewski et al. (2023); Charron et al. (2023); Batatia et al. (2023); Kovács et al. (2023); Shoghi et al. (2023); Gasteiger et al. (2022); Deng et al. (2023).
@@
MD simulations aim to accurately estimate system observables like the radial distribution function, virial stress tensor, and diffusivity coefficient. This often requires long simulation to fully explore the underlying PES . Unfortunately, MLFFs are known to produce unstable simulations, meaning that they can irreversibly enter unphysical regions of phase space (e.g., a bond breaking event in a non-reactive system at low temperature) Fu et al. (2022); Stocker et al. (2022); Vita & Schwalbe-Koda (2023); Bihani et al. (2023); Morrow et al. (2023); Wang et al. (2023b). Sampling of such regions can lead to inaccuracies in computed observables as the MLFF gradually drifts from the distribution of its training data. In extreme cases, instabilities can lead to unrecoverable simulation collapse, in which case computed observables may have high statistical error due to insufficient sampling. In either case, instability can limit the ability of MLFF-based MD simulations to investigate long-timescale phenomena like ion diffusion and protein folding, as well as rare events that may require extensive sampling to encounter. Figure 1 illustrates typical MLFF instability behaviors and provides examples of unstable structures sampled during MD simulation.
@@
Simulation instability has been shown to have an unreliable correlation with the energy and force error metrics typically used to train and evaluate MLFFs Fu et al. (2022); Stocker et al. (2022); Bihani et al. (2023). Recent works have introduced alternative simulation-based objectives, such as localized Wang et al. (2023b) or reweighted Ge et al. (2024) energy and force errors, to achieve better downstream stability. However, these approaches do not allow the MLFF to visit new configurations, limiting the distribution over which it is trained and thus the potential improvement. Alternatively, expanding the phase space coverage of the dataset can be an effective way to combat MLFF instability Fu et al. (2022); Stocker et al. (2022); Vita & Schwalbe-Koda (2023); Bihani et al. (2023); Morrow et al. (2023); Wang et al. (2023b). This is typically accomplished via active learning Smith et al. (2018); Vandermause et al. (2020); Schran et al. (2020); Lin et al. (2021); Kulichenko et al. (2023) approaches, where new atomistic configurations are selected, ab-initio QM calculations are performed, and the MLFF is retrained on the expanded dataset. Howver, these techniques rely on performing additional ab-initio calculations to expand the dataset. With MLFFs being trained on increasingly diverse datasets Shoghi et al. (2023); Batatia et al. (2023); Kovács et al. (2023) and used to simulate larger atomistic systems, the expense of these calculations may hamper the practicality of active learning workflows. This suggests the need for new training strategies or additional sources of information beyond energies and forces to train stable MLFFs.
@@
In this work, we bridge this gap by using both system observables and ab-initio QM data to improve MLFF stability. We introduce Stability-Aware Boltzmann Estimator (StABlE) Training, which uses efficient, parallelized MD simulations to rapidly explore regions of molecular phase space where the MLFF becomes unstable, followed by a targeted refinement of these regions using reference system observables. The key to efficient and numerically stable training lies in the Boltzmann Estimator, which enables end-to-end gradient-based learning without backpropagating directly through MD simulations. We also introduce a localized version of the Boltzmann Estimator, which enables targeted refinement of local instabilities. This is particularly important for stabilizing simulations of large, condensed-phase systems. StABlE Training is a self-contained, efficient process that leverages learning signals from both reference observables and existing QM data, with no reliance on performing additional QM reference calculations.
@@
We demonstrate StABlE Training on three systems and MLFF architectures: 1) simulation of aspirin with SchNet Schütt et al. (2018), 2) simulation of the Ac-Ala3-NHMe tetrapeptide with NequIP Batzner et al. (2022), and 3) simulation of an all-atom water system with GemNet-T Gasteiger et al. (2021). Relative to MLFFs trained solely on energies and forces, our StABlE-trained models produce significantly more stable MD simulations, recover observables more accurately, exhibit better generalization to unseen simulation temperatures, and outperform models trained on 50 times larger, labeled datasets. Our results suggest that utilizing both quantum-mechanical and observable-based modalities is required to fully exploit the available learning signal in reference datasets and train stable and accurate MLFFs. To our knowledge, StABlE Training is the first method that combines these data modalities to improve the stability of neural network potentials in MD simulations.
@@ 2 Preliminaries
Molecular dynamics simulation is used to evolve the positions and momenta of an atomistic system. Given a system of $N$ atoms, its state at time $t$ is defined by $\Gamma(t)=\{r(t),p(t)\}$, where $r(t),p(t)\in\mathbb{R}^{N\times 3}$ are the position and momenta of the atoms. We assume the system has a scalar-valued Hamiltonian function $\mathcal{H}:\mathbb{R}^{N\times 3}\times\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ of the form $\mathcal{H}(\Gamma)=\sum_{i=1}^{N}\frac{p^{{(i)}^{2}}}{2m^{(i)}}+U(r)\,$ where $U:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ is a potential energy function and $m^{(i)}$ and $p^{(i)}$ are the mass and momentum of atom $i$. By updating in the direction of the per-atom forces $-\frac{\partial U}{\partial r}$ using a numerical integration scheme such as Langevin dynamics Bussi & Parrinello (2007), a sequence of $K$ simulation states $\{\Gamma(t)\}_{t=0}^{K}$ is produced.
@@
Observables have been used extensively in the historical development of potentials Cornell et al. (1995); Marrink et al. (2007); Li et al. (2018), and more recently for MLFFs Thaler & Zavadlav (2021); Wang et al. (2020) in the context of condensed-phase Wang et al. (2023a) and titanium Röcken & Zavadlav (2023) systems, enhanced sampling of rare events Šípka et al. (2023), and protein folding simulation Ingraham et al. (2019); Navarro et al. (2023); Kolloff & Olsson (2023). Training with observables requires an efficient way to compute gradients through MD simulations while avoiding numerical instability and memory limitations. The approach is appealing due to the lack of reliance on expensive ab-initio quantum mechanical calculations, and the possibility of improved empirical consistency in settings where the underlying ab-initio method may be unreliable Cheetham & Seshadri (2024). However, past works do not consider MLFFs with stability problems, and dispense entirely with using QM data, thus losing the first-principles guarantees of matching an ab-initio PES.
@@ 3.1 Boltzmann Estimator
The non-trivial quantity to compute is the Jacobian, ${\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}/{\partial\theta}$. One way to estimate it is by using a chain rule expansion that corresponds to each step of the unrolled MD simulation Wang et al. (2020). The adjoint method Chen et al. (2019) can be used to limit the memory footprint at the expense of increased computation, but can still lead to numerical instability for long trajectories Wang et al. (2023a); Šípka et al. (2023).
We can avoid direct backpropagation through the simulation by noting that the equilibrium state distribution $P_{\theta}(\Gamma)$ is independent of the algorithm (i.e., MD integrator) used to sample the distribution. This is analogous to implicit differentiation techniques for differentiable optimization Amos & Kolter (2017); Gould et al. (2016); Ren et al. (2022); Blondel et al. (2022); Négiar et al. (2023), in which the solution to an optimization problem is decoupled from the numerical solver used to obtain it. We leverage the known Boltzmann form of $P_{\theta}(\Gamma)$ to construct an unbiased estimator of the Jacobian.
@@
This estimator provides an unbiased estimate of the Jacobian $\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$. A proof is provided in Supplementary Section A.1. The estimator is related to the REINFORCE trick Williams (1992) and policy gradient estimators from reinforcement learning Silver et al. (2014) when a Boltzmann state distribution is assumed. The result can also be derived using thermodynamic perturbation theory Zwanzig (1954); Thaler & Zavadlav (2021).
@@ 3.2 Stability-Aware Boltzmann Estimator (StABlE) Training
The simulation phase aims to explore the molecular phase space and pinpoint regions where the MLFF becomes unstable. We sample $R$ equilibrium states from the training dataset as initial conditions for separate MD trajectories (replicas). Using the pre-trained MLFF $U_{\theta}$, we run MD simulations on these replicas in parallel for $t$ timesteps. Due to our use of efficient vectorized GPU operations, simulating many replicas in parallel has similar computational cost to simulating a single replica. We apply a predetermined stability criterion (Supplementary Section A.5) to each replica, freezing those marked unstable at their current states. Simulations continue for the remaining replicas, with stability checks every $t$ timesteps. Once a specified fraction of replicas become unstable, we rewind all unstable replicas by $t$ timesteps. This ensures that further simulation for $t$ timesteps will trigger instability, which is then corrected in the next phase. We note that many other strategies exist to explore phase space and find unstable regions, including Diffusion Monte Carlo with fictitious masses Li et al. (2021) or uncertainty-based sampling Kulichenko et al. (2023). StABlE Training can be flexibly used with any of these strategies without impacting the learning stage (described next) of the procedure.
@@
In practice, the mapping between a sparse set of system observables and a potential energy function is non-unique Noid (2013). Consequently, learning with observables alone is underconstrained. To combat this, we regularize the observable loss function (Equation 2) with the energy and forces loss function (Equation 1). The final StABlE loss function thus becomes
@@
where $\mathcal{L}_{\text{obs}}$ and $\mathcal{L}_{\text{QM}}$ were defined in Equations 2 and 1 respectively, and the hyperparameter $\lambda$ controls the strength of the regularization. Crucially, $\mathcal{L}_{\text{QM}}$ is still only computed over the original training dataset $\mathcal{D}_{train}$, and not over new structures explored during MD simulation. Therefore, the regularization requires no additional quantum mechanical energy and forces calculations.

### v3

@@ Abstract
Machine learning force fields (MLFFs) are an attractive alternative to ab-initio methods for molecular dynamics (MD) simulations. However, they can produce unstable simulations, limiting their ability to model phenomena occurring over longer timescales and compromising the quality of estimated observables. To address these challenges, we present Stability-Aware Boltzmann Estimator (StABlE) Training, a multi-modal training procedure which leverages joint supervision from reference quantum-mechanical calculations and system observables. StABlE Training iteratively runs many MD simulations in parallel to seek out unstable regions, and corrects the instabilities via supervision with a reference observable. We achieve efficient end-to-end automatic differentiation through MD simulations using our Boltzmann Estimator, a generalization of implicit differentiation techniques to a broader class of stochastic algorithms. Unlike existing techniques based on active learning, our approach requires no additional ab-initio energy and forces calculations to correct instabilities. We demonstrate our methodology across organic molecules, tetrapeptides, and condensed phase systems, using three modern MLFF architectures. StABlE-trained models achieve significant improvements in simulation stability, data efficiency, and agreement with reference observables. Crucially, the stability improvements cannot be matched by simply reducing the simulation timestep, meaning that StABlE Training effectively allows for larger timesteps in MD simulations. By incorporating observables into the training process alongside first-principles calculations, StABlE Training can be viewed as a general semi-empirical framework applicable across MLFF architectures and systems. This makes it a powerful tool for training stable and accurate MLFFs, particularly in the absence of large reference datasets. Our code is publicly available at https://github.com/ASK-Berkeley/StABlE-Training.
@@ 1 Introduction
Molecular dynamics (MD) simulation is a staple method of computational science, enabling high-resolution spatiotemporal modeling of atomistic systems throughout biology, chemistry, and materials science (Frenkel & Smit, 2001).While the atomic forces needed for MD simulation can be obtained on-the-fly via ab-initio quantum-mechanical (QM) calculations (Car & Parrinello, 1985), this is prohibitively expensive for realistic length and time scales (Friesner, 2005). Machine learning force fields (MLFFs) have recently emerged as a promising option to serve as surrogate models for QM calculations, demonstrating the ability to capture complex many-body interactions, and in some cases transfer flexibly across chemical space (Schütt et al., 2018; Hu et al., 2021; Liu et al., 2022; Gasteiger et al., 2021; 2020; Batzner et al., 2022; Musaelian et al., 2022; Batatia et al., 2022; Schütt et al., 2021). Graph neural network (GNN)-based MLFFs trained on large ab-initio datasets are increasingly being used to model challenging and important chemical systems with favorable results (Merchant et al., 2023; Kozinsky et al., 2023; Chen & Ong, 2022; Schaarschmidt et al., 2022; Takamoto et al., 2022; Majewski et al., 2023; Charron et al., 2023; Batatia et al., 2023; Kovács et al., 2023; Shoghi et al., 2023; Gasteiger et al., 2022; Deng et al., 2023).
@@
MD simulations aim to accurately estimate system observables like the radial distribution function, virial stress tensor, and diffusivity coefficient. This often requires long simulation to fully explore the underlying PES . Unfortunately, MLFFs are known to produce unstable simulations, meaning that they can irreversibly enter unphysical regions of phase space (e.g., a bond breaking event in a non-reactive system at low temperature) (Fu et al., 2022; Stocker et al., 2022; Vita & Schwalbe-Koda, 2023; Bihani et al., 2023; Morrow et al., 2023; Wang et al., 2023b). Sampling of such regions can lead to inaccuracies in computed observables as the MLFF gradually drifts from the distribution of its training data. In extreme cases, instabilities can lead to unrecoverable simulation collapse, in which case computed observables may have high statistical error due to insufficient sampling. In either case, instability can limit the ability of MLFF-based MD simulations to investigate long-timescale phenomena like ion diffusion and protein folding, as well as rare events that may require extensive sampling to encounter. Figure 1 illustrates typical MLFF instability behaviors and provides examples of unstable structures sampled during MD simulation.
@@
Simulation instability has been shown to have an unreliable correlation with the energy and force error metrics typically used to train and evaluate MLFFs (Fu et al., 2022; Stocker et al., 2022; Bihani et al., 2023). Recent works have introduced alternative simulation-based objectives, such as localized (Wang et al., 2023b) or reweighted (Ge et al., 2024) energy and force errors, to achieve better downstream stability. However, these approaches do not allow the MLFF to visit new configurations, limiting the distribution over which it is trained and thus the potential improvement. Alternatively, expanding the phase space coverage of the dataset can be an effective way to combat MLFF instability (Fu et al., 2022; Stocker et al., 2022; Vita & Schwalbe-Koda, 2023; Bihani et al., 2023; Morrow et al., 2023; Wang et al., 2023b). This is typically accomplished via active learning (Smith et al., 2018; Vandermause et al., 2020; Schran et al., 2020; Lin et al., 2021; Kulichenko et al., 2023) approaches, where new atomistic configurations are selected, ab-initio QM calculations are performed, and the MLFF is retrained on the expanded dataset. However, these techniques rely on performing additional ab-initio calculations to expand the dataset. With MLFFs being trained on increasingly diverse datasets (Shoghi et al., 2023; Batatia et al., 2023; Kovács et al., 2023) and larger atomistic systems, the expense of these calculations may hamper the practicality of active learning workflows. This suggests the need for additional sources of information beyond energies and forces to train stable MLFFs.
@@
In this work, we bridge this gap by using both system observables and ab-initio QM data to improve MLFF stability. We introduce Stability-Aware Boltzmann Estimator (StABlE) Training, woa procedure designed to produce MLFFs that are both stable and accurate. The core idea behind StABlE is to which uses efficient, parallelized MD simulations to rapidly explore regions of molecular phase space where the MLFF becomes unstable, followed by a targeted refinement of these regions using reference system observables. The key to efficient and numerically stable training lies in the Boltzmann Estimator, which enables end-to-end gradient-based learning without backpropagating directly through MD simulations. We also introduce a localized version of the Boltzmann Estimator, which enables targeted refinement of local instabilities. This is particularly important for stabilizing simulations of large, condensed-phase systems. StABlE Training is a self-contained, efficient process that leverages learning signals from both reference observables and existing QM data, with no reliance on performing additional QM reference calculations.
@@
We demonstrate StABlE Training on three systems and MLFF architectures: 1) simulation of aspirin with SchNet (Schütt et al., 2018), 2) simulation of the Ac-Ala3-NHMe tetrapeptide with NequIP (Batzner et al., 2022), and 3) simulation of an all-atom water system with GemNet-T (Gasteiger et al., 2021). Relative to MLFFs trained solely on energies and forces, our StABlE-trained models produce significantly more stable MD simulations, recover observables more accurately, exhibit better generalization to unseen simulation temperatures, and outperform models trained on 50 times larger, labeled datasets. Our results suggest that utilizing both quantum-mechanical and observable-based modalities is required to fully exploit the available learning signal in reference datasets and train stable and accurate MLFFs. To our knowledge, StABlE Training is the first method that combines these data modalities to improve the stability of neural network potentials in MD simulations.
@@ 2 Preliminaries
Molecular dynamics simulation is used to evolve the positions and momenta of an atomistic system. Given a system of $N$ atoms, its state at time $t$ is defined by $\Gamma(t)=\{r(t),p(t)\}$, where $r(t),p(t)\in\mathbb{R}^{N\times 3}$ are the position and momenta of the atoms. We assume the system has a scalar-valued Hamiltonian function $\mathcal{H}:\mathbb{R}^{N\times 3}\times\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ of the form $\mathcal{H}(\Gamma)=\sum_{i=1}^{N}\frac{p^{{(i)}^{2}}}{2m^{(i)}}+U(r)\,$ where $U:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ is a potential energy function and $m^{(i)}$ and $p^{(i)}$ are the mass and momentum of atom $i$. By updating in the direction of the per-atom forces $-\frac{\partial U}{\partial r}$ using a numerical integration scheme such as Langevin dynamics (Bussi & Parrinello, 2007), a sequence of $K$ simulation states $\{\Gamma(t)\}_{t=0}^{K}$ is produced.
@@
Observables have been used extensively in the historical development of classical potentials (Cornell et al., 1995; Marrink et al., 2007; Li et al., 2018), and more recently are gaining traction as a complementary data source for training MLFFs (Wang et al., 2020; Thaler & Zavadlav, 2021; Fuchs et al., 2025). Observables have been used to train MLFFs in the context of condensed-phase (Wang et al., 2023a) and titanium (Röcken & Zavadlav, 2023) systems, enhanced sampling of rare events (Šípka et al., 2023), and protein folding simulations (Ingraham et al., 2019; Navarro et al., 2023; Kolloff & Olsson, 2023). Training with observables requires an efficient way to compute gradients through MD simulations while avoiding numerical instability and memory limitations. The approach is appealing due to the lack of reliance on expensive ab-initio quantum mechanical calculations, and the possibility of improved empirical consistency in settings where the underlying ab-initio method may be unreliable (Cheetham & Seshadri, 2024). However, past works do not consider MLFFs with stability problems, and dispense entirely with using QM data, thus losing the first-principles guarantees of matching an ab-initio PES.
@@ 3.1 Boltzmann Estimator
The non-trivial quantity to compute is the Jacobian, ${\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}/{\partial\theta}$. One way to estimate it is by using a chain rule expansion that corresponds to each step of the unrolled MD simulation (Wang et al., 2020). The adjoint method (Chen et al., 2019) can be used to limit the memory footprint at the expense of increased computation, but can still lead to numerical instability for long trajectories (Wang et al., 2023a; Šípka et al., 2023) (see Supplementary Section A.13 for an empirical comparison with direct and adjoint-based backpropagation).
We can avoid direct backpropagation through the simulation by noting that the equilibrium state distribution $P_{\theta}(\Gamma)$ is independent of the algorithm (i.e., MD integrator) used to sample the distribution. This is analogous to implicit differentiation techniques for differentiable optimization (Amos & Kolter, 2017; Gould et al., 2016; Ren et al., 2022; Blondel et al., 2022; Négiar et al., 2023), in which the solution to an optimization problem is decoupled from the numerical solver used to obtain it. We leverage the known Boltzmann form of $P_{\theta}(\Gamma)$ to construct an unbiased estimator of the Jacobian.
@@
This estimator provides an unbiased estimate of the Jacobian $\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$. A proof is provided in Supplementary Section A.1. The estimator is related to the REINFORCE trick (Williams, 1992) and policy gradient estimators from reinforcement learning (Silver et al., 2014) when a Boltzmann state distribution is assumed. The result can also be derived using thermodynamic perturbation theory (Zwanzig, 1954; Thaler & Zavadlav, 2021).
@@ 3.2 Stability-Aware Boltzmann Estimator (StABlE) Training
The simulation phase aims to explore the molecular phase space and pinpoint regions where the MLFF becomes unstable. We sample $R$ equilibrium states from the training dataset as initial conditions for separate MD trajectories (replicas). Using the pre-trained MLFF $U_{\theta}$, we run MD simulations on these replicas in parallel for $t$ timesteps. Due to our use of efficient vectorized GPU operations, simulating many replicas in parallel has similar computational cost to simulating a single replica. We apply a predetermined stability criterion (Supplementary Section A.5) to each replica, freezing those marked unstable at their current states. Simulations continue for the remaining replicas, with stability checks every $t$ timesteps. Once a specified fraction of replicas become unstable, we rewind all unstable replicas by $t$ timesteps. This ensures that further simulation for $t$ timesteps will trigger instability, which is then corrected in the next phase. We note that many other strategies exist to explore phase space and find unstable regions, including Diffusion Monte Carlo with fictitious masses (Li et al., 2021) or uncertainty-based sampling (Kulichenko et al., 2023). StABlE Training can be flexibly used with any of these strategies without impacting the learning stage (described next) of the procedure.
@@
In practice, the mapping between a sparse set of system observables and a potential energy function is non-unique (Noid, 2013). Consequently, learning with observables alone is underconstrained. To combat this, we regularize the observable loss function (Equation 2) with the energy and forces loss function (Equation 1). The final StABlE loss function thus becomes,
@@
where $\mathcal{L}_{\text{obs}}$ and $\mathcal{L}_{\text{QM}}$ were defined in Equations 2 and 1 respectively, and $\lambda$ controls the strength of the regularization. Crucially, $\mathcal{L}_{\text{QM}}$ is only computed over the original training dataset $\mathcal{D}_{train}$, and not over new structures explored during MD simulation. Therefore, the regularization requires no additional ab-initio calculations.